Diffusion Morning Report 09/11/08

Ted Berg tberg at stanford.edu
Thu Sep 11 06:11:02 PDT 2008


Good Morning All,
    Looking good this morning . Several toos going down because of CB 
change:Ebeam , Hitachi SEM , and 4160 SEM.  Hopefully they will be back 
up this afternoon. LTO testing and running-going down for pull and clean 
on Tuesday. Tylan nitride has intermittant mfs alarms. No other issues 
at this time.

  	  	  	  	  	 
  	Date: 	9/11/08 	THICKNESS-Change at 	  	 
Tube Name/Number 	Status 	  	  	  	 
Tylan 1 	up 	  	  	  	 
Tylan 2 	up 	  	  	  	 
Tylan 3 	up 	  	  	  	 
Tylan 4 	up 	  	  	  	 
Tylan 5-BBr3 	up 	  	  	  	 
Tylan 6-POCl3 	up 	  	  	  	 
Tylan 7-TEOS2 	up 	  	18.09/30 	  	 
Tylan 8-SiGe 	up 	  	25.93/30 	  	 
Tylan 9-Poly 	up 	logs not filled out!! 	0/30 	  	 
Tylan 10-Nitride 	yellow 	mfs alarms 	2.14/30 	  	 
Tylan 11-LTO 	yellow 	testing -running  	30.12/30 	  	 
Tylan 12-FGA 	up 	  	  	  	 
Thermco 1 	up 	  	  	  	 
Thermco 2 	up 	  	  	  	 
Thermco 3 	future 	  	  	  	 
Thermco4 	future 	  	  	  	 
Thermco Poly1 	up 	  	  	  	 
Thermco Poly2 	yellow 	non uniformity 	  	  	 
Thermco Nitride1 	up 	  	  	  	 
Thermco Nitride2 	future 	  	  	  	 
TEL Spa 	up 	  	  	  	 
Burn Box 	up 	
	   	  	 
Ebeam 	down 	breaker change 	  	  	 
ZYGO 	up 	  	  	  	 
FGA2 	up 	  	  	  	 
SiH4 for furnaces 	  	B on line 700 psi 	 changeover working well  	  	 


-------------- next part --------------
An HTML attachment was scrubbed...
URL: <http://snf.stanford.edu/pipermail/diffusion/attachments/20080911/1767a02d/attachment.html>


More information about the diffusion mailing list